# Semiformal method I. Pólya’s theorem for the complete graph $K_{\mathbb{N}}$ with complex edge weights

### v3

@@ 2. Indeed,
Let $V(x)\in\mathbb{C}[[x]]$ and $V(x)^{2}=tx^{2l}(1+t_{1}x+t_{2}x^{2}+\dots)$ where $l\in\mathbb{N}_{0}$ and $t,t_{j}\in\mathbb{C}$ (every square in $\mathbb{C}[[x]]$ has this form). Then there exists $v\in\mathbb{C}$ such that $v^{2}=t$ and
@@
Proof. Let $n\in\mathbb{N}_{0}$ and a,ba,\,b be formal variables. By the Vandermonde identity,
@@
Let $U(x),V(x)\in\mathbb{C}[[x]]$ and $U(x)=V(x)^{2}=tx^{2l}(1+t_{1}x+t_{2}x^{2}+\dots)$ where $l\in\mathbb{N}_{0}$ and $t,t_{j}\in\mathbb{C}$. If $c=|t_{1}|+|t_{2}|+\dots<1$, then $U(1)_{2}$ and $V(1)_{2}$ exist, and
@@
Proof. Clearly, $U(1)_{2}$ exists. By Lemma 2.28 there is $v\in\mathbb{C}$ such that $v^{2}=t$ and
@@
with $l$ zero summands. We show that it absolutely converges. Let $T$ be any finite subseries of $S$. Indeed, since $\big|\binom{1/2}{m}\big|\leq 1$ for every $m\in\mathbb{N}$, the sum of absolute values of summands in $T$ is at most
@@
By Proposition 2.4, $S$ has the sum $v_{0}+v_{1}+\dots=V(1)_{2}$. By item 1 of Proposition 2.25, $U(1)_{2}=\big(V(1)_{2}\big)^{2}$. $\Box$
@@
Finally, we consider convex light weights.
@@
For every convex light weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ we have
$\sum_{n=0}^{\infty}h(W(n))\cdot x^{n}=\frac{1}{1-x}$}\,.
Proof. Let $h$ be a convex light weight. We prove by induction on $n\in\mathbb{N}_{0}$ that $h(W(n))=1$. For $n=0$ this holds as $W(0)=\{\langle 1\rangle\}$. Let $n>0$. Recall that $\ell(w)$ denotes the last vertex of a walk $w$. Using Propositions 2.2 and 2.4, the convexity assumption and induction, we get
@@
$=$ ${\textstyle\sum_{w^{\prime}\in W(n-1)}h(w^{\prime})\cdot 1=1\,.}$
@@ 3 Extensions of Pólya’s theorem with $v=1$
We establish the first group of generalizations of Pólya’s theorem in Theorems 3.23.7. We consider the case $v=1$ when the vertex visited by walks coincides with their starting vertex.
We introduce some generating functions. Let $n\in\mathbb{N}_{0}$, $v\in\mathbb{N}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. Recall that $W(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, and that $W(v,n)\subset W(n)$ is the set of walks that visit $v$ in step $i>0$. We are interested in the generating function
@@
and define
@@
all in $\mathbb{C}[[x]]$. Here $W_{b}(n)$ are the walks of length $n$ in $K_{\mathbb{N}}$ that start and end at $1$, and $W_{c}(n)$ are the walks of length $n$ in $K_{\mathbb{N}}$ that start and end at $1$, and avoid $1$ between. The coefficients $a_{n}^{h}$, $\dots$, $d_{n}^{h}$ are correctly defined because they are sums of subseries of absolutely convergents series. We have $a_{0}^{h}=c_{0}^{h}=0$ and $b_{0}^{h}=d_{0}^{h}=1$.
@@
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight and generating functions $A_{h}(x)$, $B_{h}(x)$, $C_{h}(x)$ and $D_{h}(x)$ be as above. The following relations hold between them.
@@
Proof. 1. We show that $a_{n}^{h}=\sum_{j=0}^{n}c_{j}^{h}\cdot d_{n-j}^{h}$ for every $n\in\mathbb{N}_{0}$. For $n\leq 1$ it holds, $a_{n}^{h}=c_{n}^{h}=0$. For $n\geq 2$ the set $W(1,n)$ is countable and we split every walk in it at the first revisit of $1$. We get a map
@@
where $w_{1}=\langle u_{0},u_{1},\dots,u_{j}\rangle$ and $w_{2}=\langle u_{j},u_{j+1},\dots,u_{n}\rangle$. It is easy to see that the map $F$ is a bijection. $F$ is weight-preserving (WP): for every weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and value $F(w)=\langle w_{1},w_{2}\rangle$ we have $h(w)=h(w_{1})h(w_{2})$. If $X$ is a set of pairs of walks in $K_{\mathbb{N}}$, we write $h(X)$ for the sum of the series
@@
provided that it absolutely converges.
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight and $n\geq 2$. Using $F$ and above propositions, we get
@@ 2. We show that
For odd $n$ it is true because then $b^{h}_{n}=c^{h}_{n}=0$ as $W_{b}(n)=W_{c}(n)=\emptyset$, and in every inner summand at least one index $n_{i}$ is odd. For even $n\geq 2$ the set $W_{b}(n)$ is countable. We split every walk in it at the visits of $1$ and get a map
@@
defined as follows. For any walk
@@
$F(w)=\langle w_{1},\,w_{2},\,\dots,\,w_{j}\rangle\ \ (\in W_{c}(n_{1})\times W_{c}(n_{2})\times\dots\times W_{c}(n_{j}),\,n_{i}=m_{i}-m_{i-1})$
where $w_{i}=\langle u_{m_{i-1}},u_{m_{i-1}+1},\dots,u_{m_{i}}\rangle$, $i\in[j]$. It is easy to see that $F$ is a bijection and that it is weight-preserving similarly as above. If $X$ is a set of tuples of walks, we write $h(X)$ for the sum of the series
$\sum_{\langle w_{1},\,\dots,\,w_{k})\in X}h(w_{1})\dots h(w_{k})$}
@@
Let $h$ be a light weight and $n\geq 2$ be even. Using $F$ and above propositions, we get
@@
For better orientation we label the fourteen theorems by the triples
$\langle x,\,y,\,z\rangle\in\{v=1,\,v\neq 1\}\times\{\mathrm{gen},\,\mathrm{con}\}\times\{U(1)_{1},\,U(1)_{2},\,U(1)_{1}=+\infty\}\,$,
where “gen” refers to general edge weights $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and “con” to convex ones. The generating functions $A_{h}(x)$, $\dots$, $D_{h}(x)$ are as above.
@@
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. Suppose that $D_{h}(1)_{1}$, $C_{h}(1)_{1}$, $B_{h}(1)_{1}$ and $A_{h}(1)_{1}$ exist. Then $B_{h}(1)_{1}\neq 0$ and
@@
Proof. By item 1 of Proposition 3.1 we have $A_{h}(x)=C_{h}(x)D_{h}(x)$. By Proposition 2.19, $A_{h}(1)_{1}=C_{h}(1)_{1}D_{h}(1)_{1}$. By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$. By Proposition 2.20 we have $B_{h}(1)_{1}\neq 0$ and $C_{h}(1)_{1}=1-\frac{1}{B_{h}(1)_{1}}$. The stated formula follows. $\Box$
@@
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. Suppose that $D_{h}(1)_{2}$ exists and $\sum_{n\geq 1}|b_{n}^{h}|<1$. Then $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $A_{h}(1)_{2}$ exists and
@@
Proof. By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$. By item 1 of Proposition 2.27, $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $C_{h}(1)_{2}$ exists and $C_{h}(1)_{2}=1-\frac{1}{B_{h}(1)_{2}}$. By item 1 of Proposition 3.1 we have $A_{h}(x)=C_{h}(x)D_{h}(x)$. By item 1 of Proposition 2.25, $A_{h}(1)_{2}$ exists and $A_{h}(1)_{2}=C_{h}(1)_{2}D_{h}(1)_{2}$ . The stated formula follows. $\Box$
Theorem 3.4 (3: $v=1$, gen, $U(1)_{1}=+\infty$)
Suppose that $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$ is a light weight such that $D_{h}(1)_{1}=+\infty$. Then
$A_{h}(1)_{1}=+\infty\,.$
Proof. Thus $d_{n}^{h}>0$ for some $n>0$ and there is a vertex $u\in\mathbb{N}\setminus\{1\}$ with $h(\{1,u\})>0$. Then the walk $\langle 1,u,1\rangle$ shows that $c_{2}^{h}>0$. Since by item 1 of Proposition 3.1 we have $a_{n}^{h}=\sum_{j=0}^{n}c_{j}^{h}d_{n-j}^{h}$, we get $a_{n+2}^{h}\geq c_{2}^{h}d_{n}^{h}$ for every $n\in\mathbb{N}_{0}$. Hence
$\sum_{j=0}^{n}a_{j+2}^{h}\geq c_{2}^{h}\cdot\sum_{j=0}^{n}d_{j}^{h}\to+\infty,\ n\to\infty\,$,}
and $A_{h}(1)_{1}=+\infty$. $\Box$
We proceed to convex weights.
@@
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a convex light weight. Suppose that $B_{h}(1)_{1}$ and $C_{h}(1)_{1}$ exist. Then $B_{h}(1)_{1}\neq 0$ and
$\lim_{n\to\infty}{\textstyle a_{n}^{h}=1-\frac{1}{B_{h}(1)_{1}}\,.}$
Proof. By item 1 of Proposition 3.1 and by Proposition 2.31 we have $A_{h}(x)=C_{h}(x)D_{h}(x)=C_{h}(x)\frac{1}{1-x}$. By Proposition 2.15, $\lim_{n\to\infty}a_{n}^{h}=C_{h}(1)_{1}$. By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$. By Proposition 2.20 we have $B_{h}(1)_{1}\neq 0$ and $C_{h}(1)_{1}=1-\frac{1}{B_{h}(1)_{1}}$. The stated formula follows. $\Box$
@@
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a convex light weight. Suppose that $\sum_{n\geq 1}|b_{n}^{h}|<1$. Then $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$ and
$\lim_{n\to\infty}{\textstyle a_{n}^{h}=1-\frac{1}{B_{h}(1)_{2}}\,.}$
Proof. By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$. By item 1 of Proposition 2.27, $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $C_{h}(1)_{2}$ exists and $C_{h}(1)_{2}=1-\frac{1}{B_{h}(1)_{2}}$. By item 1 of Proposition 3.1 and by Proposition 2.31 we have $A_{h}(x)=C_{h}(x)D_{h}(x)=C_{h}(x)\frac{1}{1-x}$. By Proposition 2.15,
$\lim_{n\to\infty}a_{n}^{h}=C_{h}(1)_{1}=C_{h}(1)_{2}=1-\frac{1}{B_{h}(1)_{2}}\,.$}
$\Box$
Theorem 3.7 (6: $v=1$, con, $U(1)_{1}=+\infty$)
Let $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$ be a convex light weight such that $B_{h}(1)=+\infty$. Then
$\lim_{n\to\infty}a_{n}^{h}=1\,.$
Proof. By item 1 of Proposition 3.1 and Proposition 2.31 we have $A_{h}(x)=C_{h}(x)D_{h}(x)=C_{h}(x)\frac{1}{1-x}$. By Proposition 2.15, $\lim_{n\to\infty}a_{n}^{h}=C_{h}(1)_{1}$. By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$. But now by Proposition 2.31 we have $0\leq b_{n}^{h},c_{n}^{h}\leq d_{n}^{h}=1$ for every $n\in\mathbb{N}_{0}$. Hence $b_{n}^{h},c_{n}^{h}\in[0,1]$ and the generating functions $B_{h}(x)$ and $C_{h}(x)$ absolutely converge for $x\in[0,1)$. By Proposition 2.21, $C_{h}(1)_{1}=1$. The stated formula follows. $\Box$

### v6

@@ 2. Indeed, the sum
Let $V(x)\in\mathbb{C}[[x]]$ and
$V(x)^{2}=tx^{2l}(1+t_{1}x+t_{2}x^{2}+\dots)$
where $l\in\mathbb{N}_{0}$ and $t,t_{j}\in\mathbb{C}$ — every square in $\mathbb{C}[[x]]$ has this form. Then there exists a number $v\in\mathbb{C}$ such that $v^{2}=t$ and
@@
Proof. Let $n\in\mathbb{N}_{0}$ and a,ba,\,b be formal variables. By Vandermonde’s identity ([30]),
@@
Let $U(x),V(x)\in\mathbb{C}[[x]]$ and
$U(x)=V(x)^{2}=tx^{2l}(1+t_{1}x+t_{2}x^{2}+\dots)$
where $l\in\mathbb{N}_{0}$ and $t,t_{j}\in\mathbb{C}$. If $c=|t_{1}|+|t_{2}|+\dots<1$, then $U(1)_{2}$ and $V(1)_{2}$ exist, and
@@
Proof. It follows from the assumption on $c$ that $U(1)_{2}$ exists. By Lemma 2.28 there is a number $v\in\mathbb{C}$ such that $v^{2}=t$ and
@@
beginning with $l$ summands 0. We show that it absolutely converges. Let $T$ be a finite subseries of $S$. Since $\big|\binom{1/2}{m}\big|\leq 1$ for every $m\in\mathbb{N}$, the sum of absolute values of summands in $T$ is at most
@@
Thus $S$ absolutely converges. By Proposition 2.4, the series $S$ has the sum $v_{0}+v_{1}+\dots=V(1)_{2}$. By item 1 of Proposition 2.24, $U(1)_{2}=\big(V(1)_{2}\big)^{2}$. $\Box$
@@
In conclusion, we consider $\xi$-convex light weights.
@@
Let $\xi\in\mathbb{C}$. For every $\xi$-convex light weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ we have
$\sum_{n=0}^{\infty}h(W(n))x^{n}=\frac{1}{1-\xi x}\,.$}
In particular, for $\xi=0$ we have $\sum_{n=0}^{\infty}h(W(n))x^{n}=1$.
Proof. Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be light and $\xi$-convex. We prove by induction on $n\in\mathbb{N}_{0}$ that $h(W(n))=\xi^{n}$. For $n=0$ this holds as $W(0)=\{\langle 1\rangle\}$. Let $n>0$. Recall that $\ell(w)$ denotes the last vertex of a walk $w$. Using Propositions 2.2 and 2.4, the $\xi$-convexity assumption and induction, we get
@@
$=$ ${\textstyle\sum_{w^{\prime}\in W(n-1)}h(w^{\prime})\cdot\xi=\xi^{n-1}\cdot\xi=\xi^{n}\,.}$
@@ 3 Extensions of Pólya’s theorem with $v=1$
We begin with the case $v=1$ when the vertex to be visited by walks coincides with their starting vertex. In Theorems 3.23.7 we establish the first group of generalizations of Pólya’s theorem.
For that end, we need some generating functions. Let $n\in\mathbb{N}_{0}$, $v\in\mathbb{N}$, and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. Recall that $W(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, and that $W(v,n)\subset W(n)$ is the subset of walks visiting $v$ in some step $i>0$. We are interested in the generating function
@@
and define three more generating functions
@@
all in $\mathbb{C}[[x]]$. Here $W_{b}(n)$ are the walks of length $n$ in $K_{\mathbb{N}}$ that start and end at $1$, and $W_{c}(n)$ are the walks of length $n$ in $K_{\mathbb{N}}$ that start and end at $1$, and avoid $1$ between. The coefficients $a_{n}^{h}$, $\dots$, $d_{n}^{h}$ are correctly defined because they are sums of subseries of the absolutely convergent series
$d_{n}^{h}=\sum_{w\in W(n)}h(w)\,.$}
The constant coefficients are $a_{0}^{h}=c_{0}^{h}=0$ and $b_{0}^{h}=d_{0}^{h}=1$.
@@
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight and $A_{h}(x)$, $B_{h}(x)$, $C_{h}(x)$ and $D_{h}(x)$ be the above generating functions. The following relations hold.
@@
Proof. 1. We show that $a_{n}^{h}=\sum_{j=0}^{n}c_{j}^{h}\cdot d_{n-j}^{h}$ for every $n\in\mathbb{N}_{0}$. For $n\leq 1$ it holds, then $a_{n}^{h}=c_{n}^{h}=0$. For $n\geq 2$ the set $W(1,n)$ is countable and we split every walk in it at the first revisit of $1$. We get a map
@@
with $w_{1}=\langle u_{0},u_{1},\dots,u_{j}\rangle$ and $w_{2}=\langle u_{j},u_{j+1},\dots,u_{n}\rangle$. It is easy to see that $F$ is a bijection. $F$ is also weight-preserving (WP): for every weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and value $F(w)=\langle w_{1},w_{2}\rangle$, we have $h(w)=h(w_{1})h(w_{2})$. If $X$ is a set of pairs of walks in $K_{\mathbb{N}}$, we write $h(X)$ for the sum of the series
@@
if it absolutely converges.
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight and let $n\geq 2$. Using the map $F$, we get
@@ 2. We show that
For odd $n$ it is true: for odd $n$ we have $b^{h}_{n}=c^{h}_{n}=0$, and in every inner summand at least one index $n_{i}$ is odd. For even $n\geq 2$ the set $W_{b}(n)$ is countable. We split every walk in it at the visits of $1$ and get a map
@@
that is defined as follows. For any walk
@@
$F(w)=\langle w_{1},\,w_{2},\,\dots,\,w_{j}\rangle\ \ (\in W_{c}(n_{1})\times W_{c}(n_{2})\times\dots\times W_{c}(n_{j}),\,n_{i}=m_{i}-m_{i-1})\,$,
with $w_{i}=\langle u_{m_{i-1}},u_{m_{i-1}+1},\dots,u_{m_{i}}\rangle$, $i\in[j]$. It is easy to see that $F$ is a bijection and that it is weight-preserving, as in part 1. If $X$ is a set of tuples of walks, we write $h(X)$ for the sum of the series
$\sum_{\langle w_{1},\,\dots,\,w_{k})\in X}h(w_{1})\dots h(w_{k})\,$,}
@@
Let $h$ be a light weight and $n\geq 2$ be even. Using the map $F$, we get
@@
We label our fourteen generalizations of Pólya’s theorem for better orientation by the triples
$\langle x,\,y,\,z\rangle\in\{v=1,\,v\neq 1\}\times\{\mathrm{gen},\,\mathrm{con}\}\times\{U(1)_{1},\,U(1)_{2},\,U(1)_{1}=\infty\}\,$,
where “gen” refers to general edge weights $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and “con” to $\xi$-convex ones. For $U(1)_{1}$ and $U(1)_{2}$ see Definition 2.12, and for $U(1)_{1}=\infty$ see Definition 2.14. The generating functions $A_{h}(x)$, $B_{h}(x)$, $C_{h}(x)$, and $D_{h}(x)$ are defined at the beginning of this section.
@@
If $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a light weight such that $A_{h}(1)_{1}$, $B_{h}(1)_{1}$, $C_{h}(1)_{1}$ and $D_{h}(1)_{1}$ exist, then $B_{h}(1)_{1}\neq 0$ and
@@
Proof. By item 1 of Proposition 3.1 we have $A_{h}(x)=C_{h}(x)D_{h}(x)$. By Proposition 2.20, $A_{h}(1)_{1}=C_{h}(1)_{1}D_{h}(1)_{1}$. By item 2 of Proposition 3.1 we have $1-C_{h}(x)=\frac{1}{B_{h}(x)}$. By Proposition 2.21 we have $B_{h}(1)_{1}\neq 0$ and
$(1-C_{h})(1)_{1}=\frac{1}{B_{h}(1)_{1}},\,\text{ so that }\,C_{h}(1)_{1}=1-\frac{1}{B_{h}(1)_{1}}\,.$}
The stated formula follows. $\Box$
@@
If $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a light weight such that $\sum_{n\geq 1}|b_{n}^{h}|<1$ and $D_{h}(1)_{2}$ exists, then $B_{h}(1)_{2}\neq 0$, and
@@
Proof. By item 2 of Proposition 3.1 we have $1-C_{h}(x)=\frac{1}{B_{h}(x)}$. By item 1 of Proposition 2.27, $B_{h}(1)_{2}\neq 0$ and
$(1-C_{h}(1))_{2}=\frac{1}{B_{h}(1)_{2}},\,\text{ so that }\,C_{h}(1)_{2}=1-\frac{1}{B_{h}(1)_{2}}\,.$}
By item 1 of Proposition 3.1 we have $A_{h}(x)=C_{h}(x)D_{h}(x)$. By item 1 of Proposition 2.24, $A_{h}(1)_{2}=C_{h}(1)_{2}D_{h}(1)_{2}$ . The stated formula follows. $\Box$
Theorem 3.4 (3: $v=1$, gen, $U(1)_{1}=\infty$)
If $\xi\in\mathbb{C}_{1}$ and $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ is a light weight such that $(D_{h})_{1/\xi}(1)_{1}=+\infty$, then
$(A_{h})_{1/\xi}(1)_{1}=+\infty\,.$
Proof. From $(D_{h})_{1/\xi}(1)_{1}=+\infty$ it follows that $\xi^{-n}d_{n}^{h}>0$ for some $n\in\mathbb{N}$. Hence there is a vertex $u\in\mathbb{N}\setminus\{1\}$ with $\xi^{-1}h(\{1,u\})>0$. The walk $\langle 1,u,1\rangle$ shows by the subseries bound principle that $\xi^{-2}c_{2}^{h}>0$. Since by item 1 of Proposition 3.1 we have $a_{n}^{h}=\sum_{j=0}^{n}c_{j}^{h}d_{n-j}^{h}$, we get that
$\xi^{-2-n}a_{n+2}^{h}\geq\xi^{-2}c_{2}\cdot\xi^{-n}d_{n}^{h}$
for every $n\in\mathbb{N}_{0}$. Hence
$\sum_{j=0}^{n}\xi^{-2-j}a_{j+2}^{h}\geq\xi^{-2}c_{2}^{h}\cdot\sum_{j=0}^{n}\xi^{-j}d_{j}^{h}\to+\infty$}
for $n\to\infty$, and $(A_{h})_{1/\xi}(1)_{1}=+\infty$. $\Box$
We proceed to $\xi$-convex weights.
@@
1. If $\xi\in\mathbb{C}_{1}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $\xi$-convex light weight and if $(B_{h})_{1/\xi}(1)_{1}$ and $(C_{h})_{1/\xi}(1)_{1}$ exist, then $(B_{h})_{1/\xi}(1)_{1}\neq 0$ and
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h}=1-\frac{1}{(B_{h})_{1/\xi}(1)_{1}}\,.$
2. If $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a 0-convex light weight, and if $B_{h}(1)_{1}$ and $C_{h}(1)_{1}$ exist, then $B_{h}(1)_{1}\neq 0$ and
$A_{h}(1)_{1}=1-\frac{1}{B_{h}(1)_{1}}\,.$
Proof. 1. By item 1 of Proposition 3.1 and by Proposition 2.31 we have $A_{h}(x)=C_{h}(x)D_{h}(x)=C_{h}(x)\frac{1}{1-\xi x}$. By Proposition 2.15,
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h}=(C_{h})_{1/\xi}(1)_{1}\,.$
By item 2 of Proposition 3.1 we have $1-C_{h}(x)=\frac{1}{B_{h}(x)}$. By Proposition 2.21 we have $(B_{h})_{1/\xi}(1)_{1}\neq 0$ and
[…]
The stated formula follows.
2. Now, by item 1 of Proposition 3.1 and by Proposition 2.31, we have
$A_{h}(x)=C_{h}(x)$
and argue as in the previous case for $\xi=1$. $\Box$
@@
1. If $\xi\in\mathbb{C}_{1}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $\xi$-convex light weight and $\sum_{n\geq 1}\big|b_{n}^{h}\big|<1$, then $(B_{h})_{1/\xi}(1)_{2}\neq 0$ and
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h}=1-\frac{1}{(B_{h})_{1/\xi}(1)_{2}}\,.$
2. If $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a 0-convex light weight and $\sum_{n\geq 1}\big|b_{n}^{h}\big|<1$, then $B_{h}(1)_{2}\neq 0$ and
$A_{h}(1)_{2}=1-\frac{1}{B_{h}(1)_{2}}\,.$
Proof. 1. By item 2 of Proposition 3.1 we have $1-C_{h}(x)=\frac{1}{B_{h}(x)}$. By item 1 of Proposition 2.27, $(B_{h})_{1/\xi}(1)_{2}\neq 0$ and
[…]
By item 1 of Proposition 3.1 and Proposition 2.31, $A_{h}(x)=C_{h}(x)D_{h}(x)=C_{h}(x)\frac{1}{1-\xi x}$. By Proposition 2.15
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h}=(C_{h})_{1/\xi}(1)_{1}=(C_{h})_{1/\xi}(1)_{2}=1-\frac{1}{(B_{h})_{1/\xi}(1)_{2}}\,.$}
2. In the previous case we have for $\xi=1$ that $C_{h}(1)_{2}=1-\frac{1}{B_{h}(1)_{2}}$. By item 1 of Proposition 3.1 and by Proposition 2.31, $A_{h}(x)=C_{h}(x)$. The stated formula follows. $\Box$
Theorem 3.7 (6: $v=1$, con, $U(1)_{1}=\infty$)
1. If $\xi\in\mathbb{C}_{1}$, $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ is a $\xi$-convex light weight and if $(B_{h})_{1/\xi}(1)_{1}=+\infty$, then
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h}=1\,.$
2. If $\xi\in\mathbb{C}_{1}$, $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}$ is a 0-convex light weight, $B_{h},C_{h}\in H_{1}$, and if $(B_{h})_{1/\xi}(1)_{1}=\pm\infty$, then
$\lim_{x\nearrow 1}F_{(A_{h})_{1/\xi}}(x)=1\,.$
Proof. 1. By item 1 of Proposition 3.1 and by Proposition 2.31 we have
$A_{h}(x)=C_{h}(x)D_{h}(x)=C_{h}(x)\frac{1}{1-\xi x}\,.$}
By Proposition 2.15,
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h}=(C_{h})_{1/\xi}(1)_{1}\,.$
By Proposition 2.31 and the subseries bound principle,
$0\leq\xi^{-n}b_{n}^{h},\,\xi^{-n}c_{n}^{h}\leq\xi^{-n}d_{n}^{h}=1$
for every $n\in\mathbb{N}_{0}$. Hence $B_{h},C_{h}\in H_{1}$. By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$. Thus
$(C_{h})_{1/\xi}(1)_{1}=\lim_{x\nearrow 1}F_{(C_{h})_{1/\xi}}(x)={\textstyle 1-\frac{1}{\lim_{x\nearrow 1}F_{(B_{h})_{1/\xi}}(x)}=1-\frac{1}{(B_{h})_{1/\xi}(1)_{1}}=1\,,}$
where in the first and third equality we used Theorem 2.18. The stated formula follows.
2. By item 1 of Proposition 3.1 and by Proposition 2.31, we have $A_{h}(x)=C_{h}(x)$. By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$. Thus
$\lim_{x\nearrow 1}F_{(C_{h})_{1/\xi}}(x)={\textstyle 1-\frac{1}{\lim_{x\nearrow 1}F_{(B_{h})_{1/\xi}}(x)}=1-\frac{1}{(B_{h})_{1/\xi}(1)_{1}}=1-\frac{1}{\pm\infty}=1\,,}$
where in the second equality we used Theorem 2.17. The stated formula follows at once. $\Box$